OpenStudy (anonymous):

Medals and fans+testimonial!!! Factor f(x) = x^3 + x^2 – 4x – 4 = 0 completely showing all work.

OpenStudy (anonymous):

@kewlgeek555

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

factor first two terms factor last two terms

ganeshie8 (ganeshie8):

x^3 + x^2 – 4x – 4 --------- -------

ganeshie8 (ganeshie8):

wats the GCF of first two terms ?

OpenStudy (anonymous):

x^2 right?

ganeshie8 (ganeshie8):

Yes !! pull it out

ganeshie8 (ganeshie8):

x^3 + x^2 – 4x – 4 x^2(x + 1) - 4x - 4 -------

ganeshie8 (ganeshie8):

now look at last two terms. wats the GCF ?

OpenStudy (anonymous):

4

ganeshie8 (ganeshie8):

yup

ganeshie8 (ganeshie8):

x^3 + x^2 – 4x – 4 x^2(x + 1) - 4x - 4 x^2(x + 1) - 4(x + 1) ---------- --------

ganeshie8 (ganeshie8):

now look at those two groups. is there anything common ?

ganeshie8 (ganeshie8):

(x + 1) is the common right ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

so u can pull it out

ganeshie8 (ganeshie8):

x^3 + x^2 – 4x – 4 x^2(x + 1) - 4x - 4 x^2(x + 1) - 4(x + 1) (x+1)(x^2-4)

ganeshie8 (ganeshie8):

are we done ? or can we still factor it ?

OpenStudy (anonymous):

they both have x

ganeshie8 (ganeshie8):

notice that x^2-4 = x^2-2^2 and use the identity : a^2-b^2 = (a+b)(a-b)

ganeshie8 (ganeshie8):

x^3 + x^2 – 4x – 4 x^2(x + 1) - 4x - 4 x^2(x + 1) - 4(x + 1) (x+1)(x^2-4) (x+2)(x^2-2^2) (x+2)(x+2)(x-2) (x+2)^2(x-2)

ganeshie8 (ganeshie8):

now its completely factored. see if it makes more or less sense

OpenStudy (anonymous):

thank you, you put it very simply

ganeshie8 (ganeshie8):

glad it mad sense :) u wlc :)

ganeshie8 (ganeshie8):

*made